A man with a mass of 50kg rides a Ferris wheel of radius 10 meters rotating in a vertical circle at a constant speed of 3.0 m/s. The normal force on the man from the seat when the man is at its lowest point.

1
Expert's answer
2022-03-11T08:27:50-0500

The normal force reaction at lowest point is given by


"N=m(g+a)"

The centripetal acceleration



"a=\frac{v^2}{R}"

Finally


"N=m\left(g+\frac{v^2}{R}\right)""N=50\left(9.8+\frac{3.0^2}{10}\right)=535\: \rm N"